# Domain-aware priors enable vertical federated learning in data-scarce coral multi-omics Sam Victor1*

### v1

@@ Results
Domain-aware VFL significantly outperforms baseline methods
We evaluated three VFL approaches on the M. capitata thermal stress classification task using leave-one-out cross-validation across five random seeds. Domain-aware VFL achieved a mean AUROC of , substantially exceeding both the NVFlare VFL baseline () and the state-of-the-art LASER method ().
Fig 1 presents the comprehensive performance comparison. Standard NVFlare VFL performance was consistent with random chance, an expected outcome in an extreme regime (, ) where gradient updates are dominated by noise. In contrast, the proposed domain-aware VFL restored a favorable signal-to-noise ratio, enabling the model to learn stable decision boundaries. LASER, despite sophisticated label-aware representation learning, achieved only modest improvements with high variance, reflecting training instability.
@@
Paired statistical testing confirmed the superiority of Domain-aware VFL with large effect sizes (Fig 2). Domain-aware VFL significantly outperforms NVFlare VFL (, Cohen’s ) at the significance level. Comparison with LASER reveals a large effect size (Cohen’s ) but fails to reach statistical significance () solely due to LASER’s extreme variance (SD=0.215). This pattern (large effect size masked by high variance) validates that domain-aware priors provide critical stability improvements that generic VFL architectures lack.
@@
Small-sample federated learning requires resilience to initialization noise. Fig 3 illustrates the distribution of AUROC scores across random seeds. Domain-aware VFL exhibits a compact, reliable performance distribution (CV=0.036), whereas NVFlare and LASER show broad, multi-modal distributions spanning from random guessing to moderate performance. The 7-fold variance reduction relative to LASER establishes that domain-aware feature selection prevents the ”representation collapse” observed in generic methods.
@@
Crucially, this selection process is deterministic. We evaluated the stability of the feature rankings by computing the Jaccard similarity between feature sets selected across different experimental runs (Fig 5). The perfect Jaccard similarity (1.00) confirms that the method consistently identifies the same biologically relevant features regardless of random seed, enabling reproducible biomarker discovery.
@@
To confirm that the high AUROC scores reflect genuine biological signal rather than overfitting to the small sample size (), we conducted a negative control experiment using permuted labels. As shown in Fig 6, training on shuffled labels resulted in an AUROC of 0.262, significantly below random chance. This ”anti-learning” phenomenon is characteristic of high-capacity models in data-scarce regimes and confirms that the model is not merely memorizing noise; if it were, it would have achieved high training accuracy on the permuted set as well. The stark contrast between the true-label performance (0.833) and the negative control (0.262) validates the integrity of the results.
@@ Discussion
Our results propose three empirically motivated design principles for enabling vertical federated learning on ultra-high-dimensional small-sample datasets where standard VFL approaches exhibit catastrophic failure modes:
Design Principle 1: Domain-aware priors appear necessary in this regime. The magnitude of improvement, from random-chance performance (AUROC 0.50) to robust classification (AUROC 0.83), suggests that domain-aware priors play a critical role in enabling signal extraction under extreme P/N ratios. Generic VFL treats features as anonymous tokens, allowing gradient updates to be dominated by irrelevant high-dimensional noise when P/N6,000. Domain-aware feature selection via gradient saliency filters this noise before federated training, enabling convergent learning. This finding challenges conventional federated learning research, which assumes sufficient data for statistical learning and focuses on algorithmic optimization. For data-scarce specialized domains (rare diseases, ecology, proteomics), domain knowledge may become the rate-limiting factor, not algorithm sophistication [19].
Design Principle 2: Stability metrics matter more than peak performance in small-sample regimes. While our approach achieves highest mean AUROC, the 7-fold variance reduction relative to LASER (SD 0.030 vs 0.215) and 4-fold reduction relative to NVFlare (SD 0.030 vs 0.125) represent the critical practical advantage.
Design Principle 3: Interpretable feature selection mitigates gradient instability in our setting and enables domain validation. Gradient saliency-selected features provide dual benefits: (1) biological plausibility validates that performance reflects true signal, not statistical overfitting [3], and (2) perfect feature ranking stability (Jaccard=1.0) confirms robust domain knowledge extraction unaffected by random noise. Interpretability is not merely post-hoc explanation but a mechanism ensuring convergence. Biologically relevant features exhibit consistent signal-to-noise ratios, stabilizing gradient flow. Interpretability is not post-hoc explanation but a mechanism supporting convergence and domain acceptance.
@@
LASER’s failure on our dataset (AUROC 0.60±0.21, barely above random) reveals a fundamental architectural limitation of latent representation alignment methods under extreme dimensionality. LASER assumes the challenge is identifying which available features are informative versus noisy, attempting to align embeddings across modalities to extract shared signal. However, when P/N6,000 with N=13 samples, both modalities’ embeddings are dominated by high-dimensional noise, LASER effectively aligns noise with noise. Without prior dimensionality reduction, LASER’s latent space collapses to random projections varying drastically across random seeds (CV=0.358), explaining both low mean performance and catastrophic variance. This failure mode generalizes to other state-of-the-art VFL methods (PLUG, SplitFed, FedOnce) optimizing gradient communication, differential privacy, or asynchronous aggregation. These advances assume abundant training data (N10,000) where the challenge is efficiency, privacy, or scalability, not statistical feasibility. Our negative control experiment (AUROC 0.26 with permuted labels) demonstrates that high-capacity models in PN regimes exhibit ”anti-learning,” fitting inverse patterns due to gradient noise. Domain-aware feature selection transforms the learning problem from impossible (P=90,579, noise-dominated) to tractable (P=1,300, signal-enriched), enabling convergent federated training. Conversely, domain-aware machine learning for multi-omics (pathway-constrained networks, knowledge graphs, multi-task learning) typically assumes centralized data, precluding adoption under privacy constraints. Our contribution uniquely demonstrates that domain-aware priors and federated architecture are synergistically necessary for privacy-preserving collaborative learning on scarce specialized data. Neither alone suffices.
@@
The three design principles may generalize to federated learning scenarios characterized by: (1) extreme PN ratios (P/N1,000), (2) established domain knowledge (biological pathways, physical laws, expert priors), and (3) privacy/regulatory constraints preventing data centralization. These implications are speculative and require empirical validation in each domain. Potential applications encompass:
Rare disease research: Orphan diseases with N100 patients, whole-genome sequencing (P20,000 genes), multi-institutional data fragmentation, HIPAA/GDPR compliance. Pathway-based feature selection using KEGG/Reactome could enable federated GWAS where current approaches fail [19]. Microbiome studies: Gut-health associations with metagenomic sequencing (P10,000 species), small cohorts (N50), distributed collection sites. Phylogenetic priors and known host-microbe interactions guide feature selection .
Precision agriculture: Crop-disease prediction with hyperspectral imaging (P100,000 wavelengths), limited field trials (N50), competitive farm data. Spectral absorption physics and known plant stress responses inform band selection. Climate science: Paleoclimate reconstructions from multi-proxy data (ice cores, sediments, tree rings), sparse historical samples (N100), international research networks. Geophysical process models constrain feature engineering.
The critical requirement is structured domain knowledge translatable to feature priors. Our gradient saliency approach using neural network embeddings generalizes via appropriate architecture selection (CNNs for imaging, transformers for sequences, MLPs for tabular data), prioritizing knowledge-centric ML over data-centric approaches.
@@
Several scope boundaries contextualize our contributions. First, our coral-holobiont-specific biological weighting scheme (transcriptomics 1.5×, microbiome 0.5×) reflects thermal stress response pathways in M. capitata, with validation on other coral species (e.g., Acropora, Pocillopora) required to assess whether these weights generalize across coral taxa or require species-specific calibration. The selected stress biomarkers (heat shock proteins, oxidative stress enzymes, photosystem markers) represent canonical coral bleaching pathways expected to transfer across reef-building corals, but validation against independent coral stress datasets remains necessary to confirm biological interpretability. Data scarcity limits multi-species validation. Only the M. capitata thermal stress dataset provides matched multi-omics at sufficient depth for federated learning evaluation. While our architectural framework (encoder-based VFL with gradient saliency) generalizes to other PN domains with structured priors, the specific biological weights and pathway knowledge are intentionally coral-centric, designed to enable privacy-preserving collaboration among coral research laboratories rather than as a domain-agnostic solution. Second, performance variance across random seeds (SD=0.030) reflects small-sample uncertainty inherent to N=13 training samples. While domain-aware VFL achieves 7-fold variance reduction compared to LASER, prospective validation on independent coral populations from different geographic locations, stress regimes, and temporal snapshots remains necessary before operational deployment in conservation monitoring programs. The demonstrated stability enables reliable federated training, but does not eliminate the need for external validation to confirm generalization beyond the experimental cohort. Third, gradient saliency computation requires neural network forward passes and backpropagation for importance estimation, introducing one-time preprocessing overhead before federated training begins. For coral monitoring scenarios, this preprocessing occurs offline during feature set curation rather than requiring real-time computation. The 98.6% communication reduction (90,579 → 1,300 features) substantially improves bandwidth efficiency for multi-institutional coral research networks with limited connectivity, offsetting preprocessing costs in practical deployments. Fourth, leave-one-out cross-validation (LOOCV), while standard for N20 datasets, provides optimistic performance estimates compared to true holdout validation. The negative control experiment (AUROC 0.262 with permuted labels) confirms absence of data leakage, but external validation on independent coral stress experiments strengthens conclusions about biological signal versus dataset-specific patterns. Broader coral multi-omics data availability remains a field-wide challenge limiting cross-study validation.
Scope and ablation limitations
We emphasize that our study does not fully disentangle the effects of dimensionality reduction from the incorporation of biological priors. While our results demonstrate that gradient saliency–guided feature selection yields substantial gains in stability and performance relative to generic VFL baselines, we do not claim that these improvements arise exclusively from domain knowledge rather than aggressive feature pruning. Systematic ablation of surrogate models, random feature selection at equivalent dimensionality, and alternative unsupervised reduction methods (PCA, autoencoders, variance filtering) is an important direction for future work. We view the present study as evidence that unconstrained representation learning becomes ill-posed under extreme PN regimes, and that some form of structured prior, whether domain-informed or data-driven, is required for tractable learning in practice. The question of whether biological domain knowledge provides unique advantages over generic dimensionality reduction strategies remains open and merits rigorous empirical investigation in future studies.
@@
While VFL protects raw omics data by transmitting only embeddings, gradient-based privacy attacks remain theoretically possible. Advanced adversaries could potentially infer feature properties from gradient patterns or embedding distributions, though practical attacks on our 64-dimensional embeddings are non-trivial. Future work should incorporate differential privacy mechanisms (gradient perturbation, private aggregation) to provide formal privacy guarantees. The communication efficiency gains from feature selection additionally reduce attack surface by limiting transmitted information volume. Gradient saliency computation on local data introduces potential privacy leakage if surrogate models are shared. Our implementation computes feature importance locally using gradient-based saliency. Feature selection masks and importance scores are currently transmitted to coordinate selection across silos. While this enables coordinated feature selection, it introduces potential privacy leakage through importance score distributions. Future work should explore privacy-preserving alternatives such as secure multi-party computation for federated feature selection or local-only importance computation with only selection masks transmitted. Nevertheless, even feature selection patterns could theoretically reveal distributional properties, warranting further investigation of privacy-preserving feature selection protocols.
@@
Several promising research directions emerge from this work:
Temporal dynamics: Extending domain-aware VFL to longitudinal coral stress experiments (time-series omics) would enable trajectory prediction of bleaching progression. Incorporating temporal saliency importance (how feature relevance shifts over stress duration) could identify early warning biomarkers for intervention.
Multi-task learning: Simultaneously predicting multiple coral phenotypes (bleaching severity, mortality risk, recovery potential, symbiont community state) in a federated multi-task framework could improve sample efficiency by sharing representations across related tasks.
Cross-species transfer: Pre-training domain-aware VFL models on well-studied coral species (e.g., Acropora with larger datasets) and fine-tuning on rare species (Porites, Pocillopora) could leverage transfer learning to overcome per-species data scarcity.
Active learning: Integrating uncertainty quantification with federated active learning could guide optimal sample selection for future experiments, prioritizing coral specimens with maximum information gain for model improvement.
Mechanistic integration: Incorporating metabolic pathway models, gene regulatory networks, or coral-symbiont interaction dynamics as structured priors (beyond feature selection) could further constrain model space and improve interpretability.
Federated explainability: Developing privacy-preserving saliency aggregation protocols where multiple silos collaboratively compute global feature importance without sharing local data could enhance federated interpretability.

### v3

@@ Results
REEF outperforms baseline methods
We evaluated three VFL approaches plus an equal-weights ablation condition on the M. capitata thermal stress classification task using leave-one-out cross-validation across five random seeds. REEF achieved a mean AUROC of , substantially exceeding both the NVFlare VFL baseline () and the state-of-the-art LASER method (); the ablation results are reported separately in the Ablation: domain priors versus dimensionality reduction alone section.
Fig 1 presents the comprehensive performance comparison. Standard NVFlare VFL performance was consistent with random chance, an expected outcome in an extreme regime (, ) where gradient updates are dominated by noise. In contrast, REEF restored a favorable signal-to-noise ratio, enabling the model to learn stable decision boundaries. LASER, despite sophisticated label-aware representation learning, achieved only modest performance above chance with high inter-seed variance (SD=0.191), reflecting training instability in the PN regime.
| Method | Mean AUROC | SD | CV |
| REEF | 0.776 | 0.039 | 0.050 |
| Equal-weights ablation (no priors) | 0.814 | 0.090 | 0.110 |
| LASER | 0.557 | 0.191 | 0.343 |
| NVFlare VFL | 0.500 | 0.125 | 0.250 |
@@
Paired statistical testing confirmed the superiority of REEF over NVFlare VFL with a large effect size (Fig 2). REEF significantly outperforms NVFlare VFL (, Cohen’s ) at the Bonferroni-corrected threshold (, two primary comparisons). The comparison with LASER yields a large effect size (Cohen’s ) but does not reach statistical significance (), primarily because LASER exhibits high inter-seed variance (SD=0.191, CV=0.343) that inflates the standard error of the paired difference. These results confirm that REEF’s domain-aware priors provide critical stability improvements; the LASER comparison, while not statistically significant, reflects a practically meaningful numerical advantage (+0.219 AUROC, 39% relative gain) with 4.9-fold variance reduction.
| Comparison | -value | Cohen’s | Significant () |
| REEF vs. NVFlare | 0.0106 | 2.265 | Yes |
| REEF vs. LASER | 0.0995 | 1.068 | No |
| REEF vs. Equal-weights ablation | 0.405 | 0.465 | No (exploratory) |
@@
Small-sample federated learning requires resilience to initialization noise. Fig 3 illustrates the distribution of AUROC scores across random seeds. REEF exhibits a compact, reliable performance distribution (CV=0.050), whereas NVFlare and LASER show broad, multi-modal distributions spanning from random guessing to moderate performance. The 3–5-fold variance reduction relative to both baselines establishes that domain-aware feature selection prevents the ”representation collapse” observed in generic methods.
Ablation: domain priors versus dimensionality reduction alone
To disentangle the contribution of biological priors from the benefit of aggressive dimensionality reduction, we conducted an equal-weights ablation in which all domain priors were removed while holding the total feature count constant at 1,300. Specifically, we replaced the biologically motivated topk ratios (transcriptomics 0.30, proteomics 0.50, metabolomics 1.00, microbiome 0.80) with equal ratios (all 0.65, yielding exactly 325 features per omic layer), and replaced the embedding weights (1.5, 1.0, 0.8, 0.5) with uniform weights (all 1.0). All other parameters were held identical to the main domain-aware benchmark including seeds, LOOCV protocol, architecture, and hyperparameters.
The equal-weights ablation achieved a mean AUROC of (CV=0.110), compared to REEF’s (CV=0.050). The difference in mean AUROC is not statistically significant (, Cohen’s ), indicating that aggressive dimensionality reduction to 1,300 features enables above-chance learning regardless of whether features are selected by domain-informed or equal allocation. However, the ablation exhibits 2.3 higher variance (SD ratio 0.090/0.039 2.317) and a CV of 0.110 versus 0.050, confirming that biological priors are specifically responsible for the stability advantage. REEF’s domain-aware priors do not significantly improve mean performance over equal-weights selection, but they reduce performance variability by more than half.
This result refines Design Principle 1: domain-aware priors appear necessary for stable federated learning in this regime, rather than for above-chance learning per se. Any aggressive dimensionality reduction to a tractable P/N ratio enables convergent training, but biologically informed selection produces reliable, deployment-ready performance across initialization seeds.
The equal-weights condition yields an additional empirical finding: when domain priors are removed and all omic layers receive equal feature budgets and embedding weights, gradient saliency analysis reveals a clear hierarchy of intrinsic discriminative signal (Fig 4).
Proteomics emerges as the dominant omic layer in all 65 fold-seed combinations (13 LOOCV folds 5 seeds, 100% dominance frequency). Mean gradient importance scores under equal-weights conditions are: proteomics 0.091, microbiome 0.069, metabolomics 0.025, transcriptomics 0.005—proteomics importance is approximately 20 higher than transcriptomics and is stable across seeds (variance across seeds is negligible relative to the inter-omic gap). This result is consistent with proteomic approaches to coral thermal stress prediction [4] and with the broader observation that protein-level measurements integrate transcriptional and post-translational regulation, potentially capturing stress state more directly than transcript abundance. Notably, our biological prior assigned transcriptomics the highest weight (1.5) and largest relative feature budget—a choice motivated by regulatory hierarchy hypotheses that this data-driven analysis does not support for this dataset. These results suggest that proteomics-prioritized weighting schemes may improve both mean performance and stability in future iterations, and more generally demonstrate that encoder-intrinsic gradient saliency under unbiased conditions can serve as an empirical tool for prior calibration in data-scarce multi-omics settings. We note that proteomics’ considerably smaller total feature space (4,054 features versus 62,038 for transcriptomics) may contribute to higher per-feature gradient magnitudes under equal-budget conditions; disentangling architectural geometry from intrinsic biological discriminability remains an open question.
@@
Crucially, this selection process is deterministic: because importance scores are computed once from a single ranking on the fixed training set, the top-20 features are identical regardless of whether K=200, K=500, or K=1000, yielding Jaccard=1.0 across all threshold pairs. This budget-invariance is a methodological property of computing a single importance ranking rather than an empirical measure of cross-seed or cross-fold stability (see Methods for full disclosure). The practical consequence is that biomarker sets are reproducible with respect to the selection budget parameter, enabling consistent downstream interpretation.
@@
To confirm that the high AUROC scores reflect genuine biological signal rather than overfitting to the small sample size (), we conducted a negative control experiment using permuted labels. As shown in Fig 6, training on shuffled labels resulted in an AUROC of 0.357, below random chance. This sub-chance performance is characteristic of high-capacity models in data-scarce regimes and confirms that the model is not merely memorizing noise; if it were, it would have achieved high training accuracy on the permuted set as well. The contrast between the true-label performance (0.776) and the negative control (0.357) is consistent with the absence of gross data leakage; however, with , a sub-chance AUROC under permuted labels lies within normal sampling variation, so this test provides a necessary but not sufficient check against overfitting.
@@ Discussion
Our results propose three empirically motivated design principles for enabling vertical federated learning on ultra-high-dimensional small-sample datasets where standard VFL approaches exhibit severe failure modes:
Design Principle 1: Domain-aware priors appear necessary for stable learning in this regime. Generic VFL treats features as anonymous tokens, allowing gradient updates to be dominated by irrelevant high-dimensional noise when P/N6,000. Domain-aware feature selection via gradient saliency filters this noise before federated training, reducing the effective dimensionality from P=90,579 to P=1,300 and enabling convergent learning. An equal-weights ablation demonstrates that aggressive dimensionality reduction alone—without biologically informed feature allocation—also enables above-chance performance (mean AUROC 0.814), indicating that the dimensionality reduction itself is necessary for any learning to occur in this regime. However, equal-weights selection produces 2.3 higher variance than domain-aware selection (CV 0.110 vs 0.050, for mean difference), establishing that domain knowledge is specifically necessary for reliable performance rather than merely above-chance performance. For deployment in conservation monitoring, where model behavior must be predictable across experimental runs, this stability distinction is the operationally critical advantage. For data-scarce specialized domains (rare diseases, ecology, proteomics), domain knowledge may therefore function as a stability prior rather than solely a performance prior—a subtler but equally important role [17].
Design Principle 2: Stability metrics matter more than peak performance in small-sample regimes. The 3–5-fold variance reduction relative to LASER (SD 0.039 vs 0.191) and NVFlare (SD 0.039 vs 0.125) represents a critical practical advantage for deployment. The equal-weights ablation sharpens this principle: a method achieving statistically indistinguishable mean AUROC (0.814 vs 0.776, ) but 2.3 higher variance is less suitable for operational coral monitoring precisely because its performance is unpredictable across seeds. The ablation’s per-seed range spans 0.714–0.952, while REEF spans 0.714–0.833. A practitioner deploying the ablation model cannot predict whether a given run will achieve 0.714 or 0.952; a practitioner deploying REEF can rely on performance remaining within a narrow band.
Design Principle 3: Interpretable feature selection mitigates gradient instability in our setting and enables domain validation. Gradient saliency-selected features provide dual benefits: (1) biological plausibility validates that performance reflects true signal, not statistical overfitting [3], and (2) deterministic, budget-invariant feature ranking (Jaccard=1.0) enables reproducible biomarker identification; this stability is a methodological property of computing importance scores once on a fixed training set, and cross-seed or cross-fold stability remains an open question (see Methods). Biologically relevant features exhibit consistent signal-to-noise ratios, stabilizing gradient flow. Interpretability is not post-hoc explanation but a mechanism supporting convergence and domain acceptance. The proteomics dominance result (65/65 fold-seed combinations, 20 gradient importance over transcriptomics under equal-weights conditions) demonstrates this concretely: encoder-intrinsic saliency audited our own biological prior and found it incorrectly weighted, providing actionable direction for future iterations that no performance metric alone could reveal.
@@
LASER’s underperformance on our dataset (AUROC ) reveals a fundamental architectural limitation of latent representation alignment methods under extreme dimensionality. LASER assumes the challenge is identifying which available features are informative versus noisy, attempting to align embeddings across modalities to extract shared signal. However, when P/N6,000 with N=13 samples, both modalities’ embeddings are dominated by high-dimensional noise, LASER effectively aligns noise with noise. Without prior dimensionality reduction, LASER’s latent space collapses to random projections varying across random seeds (CV=0.343), explaining both modest mean performance and high variance. This failure mode generalizes to other state-of-the-art VFL methods (PLUG, SplitFed, FedOnce) optimizing gradient communication, differential privacy, or asynchronous aggregation. These advances assume abundant training data (N10,000) where the challenge is efficiency, privacy, or scalability, not statistical feasibility. The sub-chance AUROC under label permutation (0.357 for REEF, 0.238 for NVFlare) reflects the end-to-end nature of the permutation test: saliency computation in the negative control runs on encoders trained against permuted labels, selecting features that co-vary with random assignments rather than true stress signatures, and the resulting federated model operates on noise-selected features against random targets. The 2:1 ratio of true-label to permuted-label AUROC (0.776/0.357) and the stronger anti-learning in NVFlare (0.238) — which lacks domain prior structure entirely — are consistent with a mechanistic interpretation in which domain-aware priors encode genuine stress biology that conflicts directionally with random label assignments, producing systematic anti-learning rather than symmetric chance-level noise. We note that limits formal statistical discrimination between mechanistic anti-learning and sampling variation; external replication is required to confirm the effect. Domain-aware feature selection transforms the learning problem from impossible (P=90,579, noise-dominated) to tractable (P=1,300, signal-enriched), enabling convergent federated training. Conversely, domain-aware machine learning for multi-omics (pathway-constrained networks, knowledge graphs, multi-task learning) typically assumes centralized data, precluding adoption under privacy constraints. Our contribution uniquely demonstrates that domain-aware priors and federated architecture are synergistically necessary for privacy-preserving collaborative learning on scarce specialized data. Neither alone suffices.
The absence of statistical significance in the REEF versus LASER comparison () warrants direct discussion. With seed pairs, the paired -test is severely underpowered when one method produces highly variable outcomes. LASER’s inter-seed standard deviation (0.191, CV=0.343) substantially inflates the variance of paired differences, collapsing the -statistic regardless of the mean difference. A method oscillating between AUROC 0.238 and 0.738 across seeds—the observed LASER range—is itself the failure mode being characterized, not a source of statistical noise obscuring an otherwise clean comparison. The REEF versus NVFlare comparison reaches significance () under identical conditions precisely because NVFlare’s failures are consistent rather than variable. We therefore interpret the REEF versus LASER result as a demonstration of LASER’s instability rather than as evidence of insufficient advantage: the same variance that prevents significance is the variance our method eliminates.
@@
The three design principles may generalize to federated learning scenarios characterized by extreme PN ratios (P/N1,000), established domain knowledge (biological pathways, physical laws, expert priors), and privacy constraints preventing data centralization—conditions arising in rare disease genomics (GWAS with pathway priors [17]), microbiome studies (phylogenetic constraints on OTU selection), and precision agriculture (spectral physics for hyperspectral imaging). These implications are speculative and require empirical validation in each domain; the critical requirement is structured domain knowledge translatable to feature priors.
@@
Several scope boundaries contextualize our contributions. First, the biological weighting scheme (transcriptomics 1.5×, microbiome 0.5×) reflects M. capitata thermal stress pathways; validation on other coral taxa (Acropora, Pocillopora) and independent stress datasets is required to confirm whether these weights generalize or require species-specific calibration. Second, performance variance across random seeds (SD=0.039) reflects N=13 uncertainty; prospective validation on independent coral populations across geographic locations and stress regimes is necessary before operational deployment. Third, gradient saliency introduces one-time preprocessing overhead before federated training, though the 98.6% communication reduction (90,579 → 1,300 features) offsets this cost in bandwidth-limited deployments. Fourth, LOOCV provides optimistic performance estimates for small samples; the negative control (AUROC 0.357 under permuted labels) is consistent with the absence of gross data leakage but is underpowered to detect subtle overfitting with N=13, and external validation remains necessary. Regulatory frameworks (HIPAA, GDPR) governing cross-institutional data sharing provide a further motivation for privacy-preserving federated approaches [12]. Fifth, gradient saliency is computed via the same encoder architecture subsequently used for federated training; this design may favour features that align with MLP inductive biases, and alternative encoder families (attention-based, pathway-constrained networks) may yield different importance rankings, warranting caution when interpreting saliency as architecture-independent biological signal.
Ablation scope and remaining open questions
Our equal-weights ablation (Section Results) partially addresses the question of whether improvements arise from domain knowledge or aggressive dimensionality reduction per se. The result—statistically indistinguishable mean AUROC but 2.3 higher variance in the ablation—demonstrates that domain knowledge specifically contributes the stability advantage, while dimensionality reduction of any kind enables above-chance learning. This resolves the primary disentanglement concern.
Several scope boundaries remain. First, the ablation tests only one form of prior removal (equal feature budgets, uniform embedding weights). Systematic comparison against alternative dimensionality reduction strategies—PCA to 1,300 components, variance-filtered top-1,300 features, random feature selection at equivalent dimensionality—would further characterize the source of performance gains. Second, the saliency-based finding that proteomics dominates gradient importance (65/65 fold-seed combinations) under equal-weights conditions suggests that our biological prior may overweight transcriptomics relative to the data-driven signal. A proteomics-prioritized weighting scheme represents a principled next experiment. Third, the ablation tests the equal-weights condition only at K=500 (1,300 total features); whether the stability advantage of domain-aware priors persists at different feature budgets remains to be examined. Fourth, the equal-weights ablation removes both components of the biological prior simultaneously—the omic-specific feature allocation ratios and the embedding weights—which prevents attributing the observed 2.3 variance increase to either component in isolation; disentangling the independent contributions of feature selection priors versus embedding priors represents an open experimental question.
@@
While VFL protects raw omics data by transmitting only embeddings, gradient-based inference attacks remain theoretically possible on 64-dimensional embeddings, and the feature selection importance scores currently transmitted across silos introduce a secondary leakage surface. Future work should incorporate differential privacy mechanisms and explore local-only importance computation with only selection masks transmitted.
@@
The most immediate next experiment is proteomics-prioritised weighting, motivated by the saliency finding that proteomics carries the dominant intrinsic discriminative signal (65/65 fold-seed combinations) despite receiving equal feature budgets—a data-driven recalibration of the biological prior. Beyond this, extending REEF to longitudinal coral stress experiments would enable bleaching trajectory prediction, while cross-species transfer learning (pre-training on well-studied Acropora datasets, fine-tuning on rarer taxa) could leverage the framework across the limited data landscape of coral biology. Privacy-preserving federated saliency aggregation—enabling global feature importance without sharing local saliency scores—remains an open architectural challenge applicable across all PN federated domains.